A person can see clearly objects lying between 15cm and 100cm from the eye.what will be the range of his vision if he wears close fitting spectacles having a power of 0.8 D?
1
Expert's answer
2019-09-09T10:54:42-0400

"d_{min1}=15cm=0.15m; d_{max1}=100cm=1m;D_{glasses}=0.8D"

"D_{max1}=\frac{1}{d_{max1}}=\frac{1}{1}=1D"

"D_{min1}=\frac{1}{d_{min1}}=\frac{1}{0.15}=6.67D"

"D_{max2}=D_{max1}-D_{glasses}=(1-0.8)=0.2D"

"D_{min2}=D_{min1}-D_{glasses}=(6.67-0.8)=5.87D"


"d_{max2}=\frac{1}{D_{max2}}=\frac{1}{0.2}=5m=500cm"

"d_{min2}=\frac{1}{D_{min2}}=\frac{1}{5.87}=0.17m=16cm"

Answer dmin2=17cm; dmax2=500cm